# Comparing YOLO11 and YOLOv8 for instance segmentation of occluded and non-occluded immature green fruits in complex orchard environment

### v1

@@ 1 Introduction
  • Data Acquisition: To collect a consumer-grade camera-based RGB images of immature fruitlets through a robotized platform for assembling a comprehensive dataset that is precisely timed before thinning, during extensive crop care requirements.
  • Dataset Preparation: To manually prepare the occlusion and non-occlusion classes of datasets through heavy effort of annotations to test the deep learning performance.
@@
  • Performance Evaluation: Assessing the models in terms of precision and inference speed to determine their efficacy in real-world agricultural applications.
@@ 2.1 YOLOv8 Overview
Performance Metrics: YOLOv8 models are designed to deliver cutting-edge performance, with impressive metrics on benchmark datasets like COCO and ImageNet, YOLOv7, YOLOv6, YOLOv6 and so on as dipicted in Figure 3.
Figure 3: YOLO11 performance results on benchmark datasets: Performance comparison of YOLO models on the COCO dataset, showcasing YOLO11’s superior accuracy and latency against predecessors and competitors. YOLO11 achieves the highest mAP scores at varied latency levels, highlighting its efficiency in real-time applications.
@@ 2.2 YOLO11 Overview
Performance Metrics: The model exhibits superior performance metrics on benchmark datasets compared to YOLOv10, YOLOv9, YOLOv8, YOLOv7, YOLOv6, PP-YOLOE+, DAMO-YOLO, YOLOX and EfficientDet as depicted in Figure 3.
@@ 3.3 YOLO11 and YOLOv8 Training for Instance Segmentation
For YOLO11-seg and YOLOv8-seg training, both models underwent identical training protocols to ensure a balanced evaluation. The dataset comprised manually annotated images of immature green fruits in varying states of occlusions processed under a uniform configuration to facilitate consistent learning across both model architectures. Each model utilized pre-trained weights to accelerate the convergence on the specialized task of the immature green fruit segmentation, harnessing previously learned features relevant to the dataset.

### v2

@@ 1 Introduction
  • Data Acquisition: To collect consumer-grade camera-based RGB images of immature fruitlets through a robotized platform for assembling a comprehensive dataset that is precisely timed before thinning, during extensive crop care requirements.
  • Dataset Preparation: To manually prepare the occlusion and non-occlusion classes of datasets through the heavy effort of annotations to test the deep learning performance.
@@
  • Performance Evaluation: Assessing the models in terms of precision, recall, F1-scores, mAP@50, GFLOPS, number of parameters, training epochs, convolution layers and inference speed to determine their efficacy in real-world agricultural applications.
@@ 2.1 YOLOv8 Overview
Performance Metrics: YOLOv8 models are designed to deliver cutting-edge performance, with impressive metrics on benchmark datasets like COCO and ImageNet, YOLOv7, YOLOv6, YOLOv6 and so on as dipicted in Figure 3a.
@@ 2.2 YOLO11 Overview
Figure 3: YOLO11 and YOLOv8 performance results on benchmark datasets, showcasing YOLO11’s superior accuracy and latency against predecessors and competitors. YOLO11 achieves the highest mAP scores at varied latency levels, highlighting its efficiency in real-time applications:(a) Performance comparison of YOLO models on the COCO dataset, ; (b) Performance comparison between YOLO11 and YOLOv8 on the COCO dataset
@@
Performance Metrics: The model exhibits superior performance metrics on benchmark datasets compared to YOLOv10, YOLOv9, YOLOv8, YOLOv7, YOLOv6, PP-YOLOE+, DAMO-YOLO, YOLOX and EfficientDet as depicted in Figure 3a and specific comparison between YOLO11 and YOLOv8 is shown in 3b.
@@ 3.3 YOLO11 and YOLOv8 Training for Instance Segmentation
Both the YOLO11 and YOLOv8 models were trained on a workstation with an Intel Xeon® W-2155 CPU @ 3.30 GHz x20 processor, NVIDIA TITAN Xp Collector’s Edition/PCIe/SSE2 graphics card, 31.1 GiB memory, and Ubuntu 16.04 LTS 64-bit operating system. Likewise, both models underwent identical training protocols to ensure a balanced evaluation. The dataset comprised manually annotated images of immature green fruits in varying states of occlusions processed under a uniform configuration to facilitate consistent learning across both model architectures. Each model utilized pre-trained weights to accelerate the convergence on the specialized task of the immature green fruit segmentation, harnessing previously learned features relevant to the dataset.